# Fast, Accurate and Perturbative Forward Modeling of Galaxy Clustering Part II: Redshift Space

### v1

@@ 5 Conclusions
The speed-up achieved for the redshift-space modeling now enables further progress in two directions. First, the model is fast enough to be employed in field-level analyses of the redshift-space galaxy density. The higher-order information contained in the density allows to break linear-order degeneracies more efficiently than the combination of power- and bispectrum [32, 33]. We will explore the power of field-level inference to constrain $f$ and $\sigma_{8}$ independently while marginalizing over the bias coefficients. Second, given the speedup, simulation-based inference in redshift space becomes possible. The planned implementation of a survey mask will elevate the forward model to the complexity of current EFT-based power- and bispectrum analyses and facilitate SBI constraints from realistic light cones.
@@ Acknowledgments
We thank Adam Andrews for insightful feedback on this manuscript and Ivana Babić, Safak Celik, Ivana Nikolac, Minh Nguyen, Moritz Singhartinger and Beatriz Tucci for useful discussion. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311”.
@@ Appendix B Velocity estimation by the Delaunay tessellation
Figure 16: Effect of the target grid size $N_{\mathrm{G,D}}$ on the velocity power spectrum estimated by the Delaunay tessellation. The Delaunay tessellation is applied to simulations with a cut-off in the initial conditions at Λ=0.10, $\Lambda=0.10,\,0.20\,h/\mathrm{Mpc}$, respectively and $N_{\mathrm{part.}}=1536^{3}$ particles. We compare the resulting velocity power spectrum at target resolutions $N_{\mathrm{G,D}}$ (subscript “lr”) to a high-resolution reference with $N_{\mathrm{G,D}}=512$ (subscript “hr”). Dashed lines indicate expected smoothing by a top-hat kernel at the respective resolutions, and they well describe the smoothing effect of $N_{\mathrm{G,D}}$.
@@
When comparing the LEFTfield model and the reference simulations in section 2.3, we choose equal $N_{\mathrm{G,D}}=256$ such that the smoothing due to the grid sizes cancels. The particle number of simulations and forward model are $1536^{3}$ and $1024^{3}$, respectively, where we have chosen the latter for the sake of computational efficiency. The expected impact on the power spectrum from differing particle numbers is indicated in figures 3, 22, and 23, and it is smaller than the perturbative accuracy for all scenarios that we consider.
@@ C.1 Density operators
    Fourth order: $\mathrm{tr}\left[M^{(1)}\,M^{(1)}\,M^{(2)}\right]$, $\left(\mathrm{tr}\left[M^{(1)}\,M^{(1)}\right]\right)^{2}$, $\mathrm{tr}\left[M^{(1)}\,M^{(3)}\right]$, $\mathrm{tr}\left[M^{(2)}\,M^{(2)}\right]$
@@ Appendix E Sampling algorithm and convergence
$f$ $0.1$ $2.0$ $0.01$ $0.513$ $0.3-1.0$
$\sigma_{0}$ $0.0$ $10^{30}$ $0.01$ $0.3$ $0.1-2.0$
$\sigma_{\epsilon,2}$ $0.0$ $10^{30}$ $3.0$ $0.0$ $0-50$
$\sigma_{\epsilon\mu,2}$ $0.0$ $10^{30}$ $4.0$ $0.0$ $0-50$
$\beta_{\partial_{\parallel}\sigma}$ $-10^{30}$ $10^{30}$ $8.0$ $0.0$ $-50-50$
$\beta_{\partial_{\parallel}\sigma^{2}}$ $-10^{30}$ $10^{30}$ $12.0$ $0.0$
$\beta_{\partial_{\parallel}\mathrm{tr}[(M^{(1)})^{2}]}$ $-10^{30}$ $10^{30}$ $25.0$ $0.0$
$\beta_{\partial_{\parallel}M^{(1)}_{\parallel,j}\partial_{i}M^{(1),ij}}$ $-10^{30}$ $10^{30}$ $23.0$ $0.0$
@@ F.1 Accuracy of the forward model
  • The perturbative accuracy of momentum and velocity is explored in figures 1 and 3, respectively, for $\Lambda=0.20\,h/\mathrm{Mpc}$ and $z=0.5$. We extend this test to a lower cut-off, $\Lambda=0.10\,h/\mathrm{Mpc}$, and higher redshifts $z=1.0$ in figures 22 and 23.
@@ F.2 Parameter inference at fixed initial conditions
We show additional parameter recovery tests in figures 26 to 31 and comment on them in section 4.2.

### v2

@@ 5 Conclusions
The speed-up achieved for the redshift-space modeling now enables further progress in two directions. First, the model is fast enough to be employed in field-level analyses of the redshift-space galaxy density. The higher-order information contained in the density allows to break linear-order degeneracies more efficiently than the combination of power- and bispectrum [32, 33]. This is illustrated for the case of fixed initial conditions in figure 14. We plan to explore the full potential of field-level analyses to put independent constraints on $f$ and $\sigma_{8}$ while marginalizing over bias and initial conditions in the future. Second, given the speedup, simulation-based inference in redshift space becomes possible. The planned implementation of a survey mask will elevate the forward model to the complexity of current EFT-based power- and bispectrum analyses and facilitate SBI constraints from realistic light cones.
@@ Acknowledgments
We thank Adam Andrews for insightful feedback on this manuscript and Ivana Babić, Safak Celik, Ivana Nikolac, Minh Nguyen, Moritz Singhartinger and Beatriz Tucci for useful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311”.
@@ Appendix B Velocity estimation by the Delaunay tessellation
Figure 17: Effect of the target grid size $N_{\mathrm{G,D}}$ on the velocity power spectrum estimated by the Delaunay tessellation. The Delaunay tessellation is applied to simulations with a cut-off in the initial conditions at $\Lambda=0.10,\,0.20\,h/\mathrm{Mpc}$, respectively and $N_{\mathrm{part.}}=1536^{3}$ particles. We compare the resulting velocity power spectrum at target resolutions $N_{\mathrm{G,D}}$ (subscript “lr”) to a high-resolution reference with $N_{\mathrm{G,D}}=512$ (subscript “hr”). Dashed lines indicate expected smoothing by a top-hat kernel at the respective resolutions, and they well describe the smoothing effect of $N_{\mathrm{G,D}}$.
@@
When comparing the LEFTfield model and the reference simulations in section 2.3, we choose equal $N_{\mathrm{G,D}}=256$ such that the smoothing due to the grid sizes cancels. The particle number of simulations and forward model are $1536^{3}$ and $1024^{3}$, respectively, where we have chosen the latter for the sake of computational efficiency. The expected impact on the power spectrum from differing particle numbers is indicated in figures 3, 23, and 24, and it is smaller than the perturbative accuracy for all scenarios that we consider.
@@ C.1 Density operators
    Fourth order: $\mathrm{tr}\left[M^{(1)}\,M^{(1)}\,M^{(2)}\right]$, $\left(\mathrm{tr}\left[M^{(1)}\,M^{(1)}\right]\right)^{2}$, $\mathrm{tr}\left[M^{(1)}\,M^{(3)}\right]$, $\mathrm{tr}\left[M^{(2)}\,M^{(2)}\right]$ .
@@ Appendix E Sampling algorithm and convergence
$f$ $0.1$ $2.0$ $0.01$ $0.513$ 0.31.00.3-1.0
$\sigma_{0}$ $0.0$ $10^{30}$ $0.01$ $0.3$ 0.12.00.1-2.0
$\sigma_{\epsilon,2}$ $0.0$ $10^{30}$ $3.0$ $0.0$ 0500-50
$\sigma_{\epsilon\mu,2}$ $0.0$ $10^{30}$ $4.0$ $0.0$ 0500-50
$\beta_{\partial_{\parallel}\sigma}$ $10^{30}$ $10^{30}$ $8.0$ $0.0$ 5050-50-50
$\beta_{\partial_{\parallel}\sigma^{2}}$ $10^{30}$ $10^{30}$ $12.0$ $0.0$
$\beta_{\partial_{\parallel}\mathrm{tr}[(M^{(1)})^{2}]}$ $10^{30}$ $10^{30}$ $25.0$ $0.0$
$\beta_{\partial_{\parallel}M^{(1)}_{\parallel,j}\partial_{i}M^{(1),ij}}$ $10^{30}$ $10^{30}$ $23.0$ $0.0$
@@ F.1 Accuracy of the forward model
  • The perturbative accuracy of momentum and velocity is explored in figures 1 and 3, respectively, for $\Lambda=0.20\,h/\mathrm{Mpc}$ and $z=0.5$. We extend this test to a lower cut-off, $\Lambda=0.10\,h/\mathrm{Mpc}$, and higher redshifts $z=1.0$ in figures 23 and 24.
@@ F.2 Parameter inference at fixed initial conditions
We show additional parameter recovery tests in figures 27 to 32 and comment on them in section 4.2.